# Sparse Auto-Encoder Interprets Linguistic Features in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) excel in tasks that require complex linguistic abilities, such as reference disambiguation and metaphor recognition/generation. Although LLMs possess impressive capabilities, their internal mechanisms for processing and representing linguistic knowledge remain largely opaque. Previous work on linguistic mechanisms has been limited by coarse granularity, insufficient causal analysis, and a narrow focus. In this study, we present a systematic and comprehensive causal investigation using sparse auto-encoders (SAEs). We extract a wide range of linguistic features from six dimensions: phonetics, phonology, morphology, syntax, semantics, and pragmatics. We extract, evaluate, and intervene on these features by constructing minimal contrast datasets and counterfactual sentence datasets. We introduce two indices—Feature Representation Confidence (FRC) and Feature Intervention Confidence (FIC)—to measure the ability of linguistic features to capture and control linguistic phenomena. Our results reveal inherent representations of linguistic knowledge in LLMs and demonstrate the potential for controlling model outputs. This work provides strong evidence that LLMs possess genuine linguistic knowledge and lays the foundation for more interpretable and controllable language modeling in future research.
Sparse Auto-Encoder Interprets Linguistic Features
in Large Language Models
Yi Jing1, Zijun Yao1, Lingxu Ran1, Hongzhu Guo2, Xiaozhi Wang1, Lei Hou1, Juanzi Li1 1Department of Computer Science and Technology, Tsinghua University 2Department of Chinese Language and Literature, Peking University jingy22@mails.tsinghua.edu.cn, lijuanzi@tsinghua.edu.cn
@@ 1 Introduction
Large language models (LLMs) exhibit excellent performance in solving tasks that require different levels of linguistic competence, such as dependency parsing Lin et al. (2022); Roy et al. (2023), reference disambiguation Iyer et al. (2023) and metaphor interpretation Wachowiak and Gromann (2023); Yerukola et al. (2024); Tian et al. (2024). While their linguistic capabilities are largely attributed to the emergence of abilities from large-scale pre-training and model size Manning et al. (2020); Allen-Zhu and Li (2023); Mahowald et al. (2024), the underlying mechanisms by which LLMs process these linguistic structures remain under-explored Saba (2023). Thus, we aim to interpret the linguistic mechanisms of LLMs by addressing the following question: Can we identify minimal components within LLMs that are responsible for distinct linguistic processing capabilities?
Figure 1: After a sentence is input into the model, its hidden states are encoded by the SAE into a sparse feature distribution. Across multiple layers, we can identify base vectors that are significantly activated and associated with the sentence’s linguistic features.
Previous attempts at interpreting the linguistic mechanisms of LLMs usually involve instructing them with expert-designed prompts, aiming to explain how these models generate particular outputs. Yin and Neubig (2022). Nevertheless, such behavior-based methods do not provide model-structure-level mechanism interpretation. Most recent works turn to establish the connection between specific linguistic capabilities of LLMs and their interior structure, such as hidden states Katz and Belinkov (2023), attention heads Wu et al. (2020), and activated neurons Sajjad et al. (2022); Huang et al. (2023). However, these approaches mostly suffer from the following challenges:
Coarse Interpretation Granularity. Linguistic mechanism interpretation aims to find the atomic linguistic structure in LLMs. However, even neurons are the most fine-grained native components of LLMs, they are observed to be activated by multiple different conditions, a phenomena that are termed as poly-semanticity Yan et al. (2024). Thus, it is necessary to extract more fine-grained structures from LLMs to interpret their linguistic mechanism.
Insufficient Causal Analysis. Current linguistic mechanism interpretations successfully identify relevant inner structure of LLMs, with activation patching Hanna et al. (2024); Nanda et al. (2023) as the most typical methodology. However, it is still challenging to verify causal relationships between linguistic abilities and corresponding internal structures, which is a prerequisite for effectively steering model behavior through interventions on corresponding mechanisms.
To address these challenges, we propose to utilize sparse auto-encoder (SAE) for interpreting linguistic mechanisms of LLMs, a framework dubbed SAELing. SAE learns a projection matrix which decomposes the hidden states of LLMs into an extremely high-dimensional feature space with sparse activation constraint, where each dimension is expected to represent a single meaning, as Figure 1 shows. Building on this, SAELing comprises two components: (1) SAELing designs sparse feature analysis for un-interpreted features that are extracted by SAE, which provides a fine-grained linguistic mechanism interpretation for LLMs; (2) SAELing proposes to manipulate the LLMs via intervening on the features with desired interpretation, which verifies the causal relationship between the features and their interpretations. It also potentially paves way to steer the linguistic behavior of LLMs.
In particular, we first establishes a hierarchical linguistic framework with annotated corpora. The framework classifies the linguistic features into six categories, including phonetics, phonology, morphology, syntax, semantics, and pragmatics. These linguistic features are wildly observed linguistic abilities, thus guarantee the feasibility to interpret their mechanisms. To interpret sparse features in SAEs, we construct minimal pairs and counterfactual sentences for each sentence in our dataset. We also introduce a causal analysis method that intervenes on specific linguistic features via the SAE and uses an LLM as a judge to assess the intervention effect. Furthermore, we present two causality evaluation metrics: the Feature Representation Confidence (FRC) score and Feature Intervention Confidence (FIC) score, which measure a feature’s ability to identify the corresponding linguistic phenomenon in the input and its ability to regulate the model output to generate the phenomenon, respectively.
We conduct a series of experiments on Llama-3.1-8B Grattafiori et al. (2024). Our experiment results show that SAELing effectively identifies key features for linguistic competence. SAELing also provides a robust way to steer LLMs by intervening on the found linguistic features.
@@ 2 Related Works
Linguistic mechanism interpretation has been a ever-chasing goal since the emergence of LLMs. We review linguistic capability evaluation for LLMs and corresponding mechanistic interpretation works. We will also introduce the basic concepts for sparse auto-encoder.
Linguistic Features in LLMs. LLMs are shown to be equipped with diverse linguistic features. Morphological studies find inflectional and derivational phenomena along with word-formation processes in LLMs Rambelli et al. (2024); Weissweiler et al. (2023). Syntactic evaluations include canonical constructions, e.g., genitives and object-complement structures (Gauthier et al., 2020; Zhang et al., 2023; Arora et al., 2024), and cross-linguistic tests Mueller et al. (2020). Semantic investigations address metaphor comprehension Tian et al. (2024); Wachowiak and Gromann (2023); Stowe et al. (2021); He et al. (2022); Liu et al. (2022), deep semantic analysis Chen et al. (2024), and output consistency Raj et al. (2023). Pragmatic benchmarks examine the interpretation of contextual cues Sileo et al. (2022); Wu et al. (2024).
Linguistic Mechanism Interpretation. LLMs excel in most of the above tasks, which spurs growing interest in explaining their linguistic capabilities. At the behavioral explanation level, methods include feature attribution, contrastive explanation Yin and Neubig (2022), surrogate model explanation, and self-explanation. At the hidden-layer explanation level, approaches comprise analyses of attention heads Wu et al. (2020), probing tasks Hahn and Baroni (2019); Arora et al. (2024); He et al. (2022), and correlation studies Liu et al. (2024) of hidden-layer activation patterns. At the neuron explanation level, research has primarily focused on analyzing the activations of linguistically relevant neurons Tang et al. (2024).
@@
where the SAE is parameterized by $\mathbf{W}_{e}\in\mathbb{R}^{(r\times d)\times d},\mathbf{b}_{e}\in\mathbb{R}^{(r\times d)}$. $r$ is the expansion ratio, defined as the factor by which the hidden state dimension is expanded. Commonly used sparse constraint include TopK Gao et al. (2024) and JupeReLU Rajamanoharan et al. (2024) functions. As each dimension of the sparse activation in $\mathbf{f}$ corresponds to a base vector in $\mathbf{W}_{e}$, this paper uses base vector to denote features extracted by SAE.
@@ 3 Methodology
SAELing consists of three key components. (1) A hierarchical linguistic structure with supporting corpora for linguistic mechanism analysis; (2) Linguistic feature analysis for interpreting SAE extracted features; and (3) Linguistic feature intervention for causal analysis and LLM steering.
Figure 2: The overall framework of SAELing. We propose a large-model linguistic mechanism framework encompassing six dimensions and select classical features from these dimensions for experimentation. The experimental workflow is as follows: (1) Construct minimal contrast and counterfactual datasets; (2) Extract features and evaluate their relevance by analyzing the activation values of base vectors on the datasets; (3) Intervene in the model output by modifying activation values and assess causality using an LLM as a judge.
3.1 Linguistic Structure
Hierarchical Linguistic Structure.
To systematically interpret the language capabilities of large models, we adopt a six level structure based on theoretical linguistics Fromkin et al. (2017): phonetics, phonology, morphology, syntax, semantics, and pragmatics. The structure follows a logical progression from the external, physical realization of sound to the internal, contextual understanding of meaning. Each linguistic capability contains several concrete linguistic features, e.g., semantics level includes metaphor, simile, etc. We provide the exact definition for each linguistic capability in Appendix A.1
@@ 3.1 Linguistic Structure
The sparse feature activation distribution of SAE is closely related to the conditions under which their corresponding linguistic features hold in linguistic knowledge. To find the linguistic features and evaluate its dominance, we propose a method to construct the dataset and analyze feature activation frequencies.
For each linguistic feature, we first construct a set of sentences that significantly align with the desired feature. The feature activation representing this linguistic feature in SAE’s hidden space will be significantly activated on these sentences. However, this is not enough to accurately identify them, as there are some background noise vectors that are activated on all sentences in the dataset and interfere with our judgment. We need to include a control group without the feature in the constructed sentences.
We introduce two types of control groups: minimal pairs and counterfactual sentences. Minimal pairs are constructed by changing only the part of a sentence that corresponds to a particular linguistic feature, while keeping all other parts unchanged. However, this approach often results in syntactically incorrect sentences.
To overcome this limitation, we also construct fully grammatically correct control groups, called counterfactual sentences, which differ from the original sentence only in terms of its linguistic features. Detailed dataset construction procedures are provided in Appendix A.2.
3.2 Feature Analysis
We propose a causal probability approach to evaluate the relationship between extracted linguistic features and their activation on sentences containing those features.
For a given feature $x$, we define two key probabilities. The Probability of Necessity (PN) quantifies how necessary the feature is for the activation of a corresponding base vector, while the Probability of Sufficiency (PS) measures the likelihood that introducing the feature triggers activation. These probabilities are then combined into a Feature Representation Confidence (FRC) score, which assesses both the representational capacity of the SAE latent space and the discriminative ability of the feature to identify the corresponding linguistic phenomenon.
During feature analysis, we calculate the FRS on both the minimal contrast dataset and the counterfactual dataset, then average the results. This average more accurately reflects the ability of the base vectors to represent the linguistic features. Detailed definitions and calculation methods are provided in Appendix D.1.
@@ 3.3 Feature Intervention
When we modify the values of SAE’s activation during forward propagation, we expect that such targeted interventions will influence the model’s behavior. However, our experiments show that altering only a small subset of features may not significantly impact the output—likely because linguistic phenomena are represented by multiple features across various layers. To assess the true impact of these interventions, we use a large language model as a judge. For each linguistic feature, we conduct both ablation and enhancement experiments. In the ablation experiment, we set the target feature’s activation to 00, and in the enhancement experiment, we set it to $10$. In both cases, we also perform baseline experiments by randomly selecting 25 base vectors from the same layer.
@@
Let $P_{abl}^{T}$ and $P_{abl}^{B}$ denote the success probabilities (i.e., the probability that the intended change in the linguistic phenomenon is observed) for the targeted and baseline ablation experiments. The normalized ablation effect is then defined as
$E_{abl}$ $=P_{abl}^{T}-P_{abl}^{B}$
$=\frac{P(Y=0\mid I_{abl}^{T})-P(Y=0\mid I_{abl}^{B})}{P(Y=0\mid I_{abl}^{T})}.$
Similarly, let $P_{enh}^{T}$ and $P_{enh}^{B}$ be the success probabilities for the targeted and baseline enhancement experiments, with $Y=1$ indicating the presence of the phenomenon. The normalized enhancement effect is given by
$E_{enh}$ $=P_{enh}^{T}-P_{enh}^{B}$
$=\frac{P(Y=1\mid I_{enh}^{T})-P(Y=1\mid I_{enh}^{B})}{1-P(Y=1\mid I_{enh}^{B})}.$
@@
When calculating FIC, if one or both of the $E$ values are negative, we incorporate a penalty coefficient $w$ to reflect the weakened or lost causality in such cases. This FIC score provides a balanced measure of how effectively targeted interventions, as opposed to random ones, influence the model’s output with respect to specific linguistic features. The details for FIC are shown in Appendix D.2.

### v2

# LinguaLens: Towards Interpreting Linguistic Mechanisms of Large Language Models via Sparse Auto-Encoder

@@ Abstract
Large language models (LLMs) demonstrate exceptional performance on tasks requiring complex linguistic abilities, such as reference disambiguation and metaphor recognition/generation. Although LLMs possess impressive capabilities, their internal mechanisms for processing and representing linguistic knowledge remain largely opaque. Prior research on linguistic mechanisms is limited by coarse granularity, limited analysis scale, and narrow focus. In this study, we propose LinguaLens, a systematic and comprehensive framework for analyzing the linguistic mechanisms of large language models, based on Sparse Auto-Encoders (SAEs). We extract a broad set of Chinese and English linguistic features across four dimensions—morphology, syntax, semantics, and pragmatics. By employing counterfactual methods, we construct a large-scale counterfactual dataset of linguistic features for mechanism analysis. Our findings reveal intrinsic representations of linguistic knowledge in LLMs, uncover patterns of cross-layer and cross-lingual distribution, and demonstrate the potential to control model outputs. This work provides a systematic suite of resources and methods for studying linguistic mechanisms, offers strong evidence that LLMs possess genuine linguistic knowledge, and lays the foundation for more interpretable and controllable language modeling in future research.
LinguaLens: Towards Interpreting Linguistic Mechanisms
of Large Language Models via Sparse Auto-Encoder
Yi Jing, Zijun Yao, Hongzhu Guo, Lingxu Ran, Xiaozhi Wang, Lei Hou, Juanzi Li [^4] Corresponding author. DCST, BNRist; KIRC, Institute for Artificial Intelligence, Tsinghua University, China Department of Chinese Language and Literature, Peking University, China jingy22@mails.tsinghua.edu.cn, lijuanzi@tsinghua.edu.cn
[…]
@@ 1 Introduction
Large language models (LLMs) demonstrate strong performance on tasks requiring different levels of linguistic competence, such as dependency parsing Lin et al. (2022); Roy et al. (2023), reference disambiguation Iyer et al. (2023), and metaphor interpretation Wachowiak and Gromann (2023); Yerukola et al. (2024); Tian et al. (2024).
Figure 1: The main linguistic features activated at different layers are observed when example sentences are input to the model. Through a Sparse Auto-Encoder, each layer’s activation values are mapped into a sparse space and the basis vectors corresponding to predefined linguistic features are extracted. According to the results, the model’s 32 layers are divided into four stages, in order: Morphology and Core Syntax, Complex Syntactic Constructions, Pragmatic Functions, and Deep Semantics and Rhetoric.
Although their linguistic abilities are often attributed to emergent capabilities from large-scale pretraining and model scale Manning et al. (2020); Allen-Zhu and Li (2023); Mahowald et al. (2024), the underlying mechanisms by which LLMs process these linguistic structures remain underexplored and lack systematic explanation Saba (2023). Therefore, our goal is to interpret the linguistic mechanisms of LLMs by addressing the following questions: (1) Can we identify the minimal components within an LLM responsible for specific linguistic processing abilities? (2) Can we comprehensively model the internal linguistic functionalities of the model?
Prior attempts to explain LLM linguistic mechanisms typically rely on expert-designed prompts that ask the model to elucidate its generation process Yin and Neubig (2022). However, such behavior-based approaches do not provide structure-level mechanistic insights. More recent work seeks to link specific linguistic capabilities to internal structures—such as hidden states Katz and Belinkov (2023), attention heads Wu et al. (2020), and activated neurons Sajjad et al. (2022); Huang et al. (2023)—but they face two main challenges:
Coarse interpretive granularity. Mechanistic interpretation aims to uncover atomic linguistic structures within LLMs. Yet even neurons—the finest native components—exhibit poly-semantic activations, responding to multiple conditions Yan et al. (2024). This necessitates extracting finer-grained structures to truly interpret linguistic mechanisms.
Limited analysis scale. Existing studies focus on one or a few linguistic features, often within a single subfield (e.g., syntax or semantics), neglecting large-scale, systematic analysis across diverse linguistic phenomena. A scalable, automated framework is needed to interpret language mechanisms comprehensively.
To address these challenges, we propose LinguaLens, a framework that utilizes a sparse auto-encoder (SAE) to interpret LLM linguistic mechanisms. The SAE learns a projection matrix that decomposes LLM hidden states into an extremely high-dimensional feature space under a sparsity constraint, where each dimension captures a single semantic concept (Figure 1). LinguaLens comprises three modules: 1. Construction of a large-scale, multilingual, counterfactual linguistic dataset to support systematic discovery of linguistic structures; 2. Sparse feature analysis to interpret the SAE-extracted features, providing fine-grained and comprehensive mechanistic insights; 3. Feature intervention, manipulating LLM behavior via targeted interventions on interpretable features to verify causal relationships and enable controlled steering of language behavior.
Specifically, we first build a large-scale hierarchical counterfactual linguistic dataset with annotated corpora, categorizing features into morphology, syntax, semantics, and pragmatics. These widely studied linguistic abilities ensure the feasibility of interpretability. We automate feature extraction via SAE activation analysis and an LLM-based agent, and introduce a causal analysis method that intervenes on SAE base vectors with an LLM judge to evaluate effects. Building on this, we analyze cross-layer function distribution and cross-lingual representation patterns differences of linguistic features.
We conduct extensive experiments on Llama-3.1-8B Grattafiori et al. (2024). Our results demonstrate that LinguaLens can effectively identify linguistic competence features at scale, laying the groundwork for further systematic analysis.
@@ 2 Related Works
Linguistic mechanism interpretation has been a ever-chasing goal since the emergence of LLMs. Researchers build linguistic datasets to evaluate the linguistic capability and to interpret linguistic mechanisms. We review linguistic datasets for LLMs and corresponding mechanistic interpretation works. We will also introduce the basic concepts for sparse auto-encoder.
Linguistic Datasets for LLMs. Previous studies have introduced numerous linguistic datasets for large-model research, which can be divided into two main categories. The first comprises minimal-pair challenge sets—such as BLiMP Warstadt et al. (2020), CLiMP Xiang et al. (2021), and SyntaxGym Gauthier et al. (2020)—that use acceptability judgments to evaluate morphosyntactic competence. The second consists of counterfactual or contrastive corpora—including CAD Sen et al. (2022), Contrast Sets Gardner et al. (2020), and Polyjuice Wu et al. (2021)—that assess model by generating factual/counterfactual pairs. These resources focus primarily on syntactic analysis and performance evaluation, and are not suited for systematic investigation of models’ internal linguistic representations.
Linguistic Mechanism Interpretation. Previous work has employed a variety of methods to study linguistic mechanisms in large language models, including attention head analysis What Does BERT Look at? An Analysis of BERT’s Attention (2019), probing classifiers Belinkov (2022); He et al. (2024), causal intervention techniques Finlayson et al. (2021); Hao and Linzen (2023), and neuron-level analyses Sajjad et al. (2022). However, these approaches have not been applied in a unified, large-scale framework to systematically chart models’ full range of linguistic capabilities.
@@
where the SAE is parameterized by $\mathbf{W}_{e}\in\mathbb{R}^{(r\times d)\times d},\mathbf{b}_{e}\in\mathbb{R}^{(r\times d)}$. $r$ is the expansion ratio, defined as the factor by which the hidden state dimension is expanded. Commonly used sparse constraint include TopK Gao et al. (2024) and JumpReLU Rajamanoharan et al. (2024) functions. As each dimension of the sparse activation in $\mathbf{f}$ corresponds to a base vector in $\mathbf{W}_{e}$, this paper uses base vector to denote features extracted by SAE.
@@ 3 Methodology
LinguaLens consists of three key components. (1) A multi-level counterfactual dataset of linguistic features supporting systematic linguistic mechanism analysis; (2) An SAE-based linguistic feature extraction method leveraging LLM agents and correlation analysis; (3) A Linguistic feature intervention method for causality validation and LLM steering.
Figure 2: The overall framework of LinguaLens. We propose a framework for the linguistic mechanisms of large-scale models that encompasses four dimensions of theoretical linguistics and a cross-lingual analysis of both Chinese and English. The experimental workflow is as follows: (1) Construct counterfactual datasets; (2) Extract features by analyzing the activation values of base vectors on the datasets; (3) Intervene in the model output by modifying activation values and assess causality using an LLM as a judge.
3.1 Linguistic Dataset
Counterfactual Methods.
Let the presence of the target linguistic phenomenon be denoted by $T\in\{0,1\}$. For every sentence $s^{+}$ with $T=1$, define the activation of SAE base vector $k$ as $a^{(1)}_{k}=a_{k}(s^{+})$. A counterfactual sentence $s^{-}$ is produced through a minimal edit that deletes or substitutes the trigger while preserving semantic content, yielding the activation $a^{(0)}_{k}=a_{k}(s^{-})$. The individual latent effect is therefore
$\tau_{k}(s)=a^{(1)}_{k}-a^{(0)}_{k}$.
Aggregating $\tau_{k}$ across all paired sentences produces
$\operatorname{EALE}_{k}=\frac{1}{N}\sum_{i=1}^{N}\tau_{k}(s_{i}),$
which can rank base vectors by their sensitivity to the specified phenomenon.
Each $s^{-}$ must satisfy three constraints:
  1. (a)
    Minimal edit: modify only the smallest unit that realises the phenomenon (e.g. replace is eaten with eats to remove passivisation).
  2. (b)
    Semantic preservation: retain propositional content, argument structure, and discourse context so that the sentence remains truth-conditionally equivalent.
@@ 3.1 Linguistic Dataset
We construct a counterfactual dataset named LinguaLens-Data, which covers multiple linguistic domains to encompass a wide range of linguistic knowledge and functions. We select a total of $145$ linguistic features from textbooks in morphology, syntax, semantics, and pragmatics, including both English and Chinese features. For each feature, we create 50 sentences that explicitly contain the target phenomenon and apply a counterfactual minimal-editing approach to generate corresponding counterfactual sentences. Each linguistic feature is annotated with its associated linguistic domain, acknowledging that some features may lie at the interface of multiple domains. This dataset provides a foundation for future systematic studies on how specific linguistic features are represented within model internals.
3.2 Feature Extraction
Building on the counterfactual framework, we treat each paired sentence $(s^{+},s^{-})$ as a mini-experiment that perturbs only the target phenomenon $T$. Let $\theta_{k}$ be a layer-specific activation threshold (the median of $a_{k}$ on the full corpus) and define the binary trigger
$Z_{k}(s)=\mathbb{I}\!\bigl{[}a_{k}(s)\geq\theta_{k}\bigr{]}.$
Probability of Sufficiency (PS).
For base vector $k$, the probability that adding the phenomenon turns the vector “on” is
$\operatorname{PS}_{k}=\Pr\!\bigl{[}\,Z_{k}^{(1)}=1\mid Z_{k}^{(0)}=0\,\bigr{]},$
where $Z_{k}^{(1)}$ and $Z_{k}^{(0)}$ are measured on $s^{+}$ and $s^{-}$, respectively.
Probability of Necessity (PN).
Conversely, the probability that the vector would switch off if the phenomenon were removed is
$\operatorname{PN}_{k}=\Pr\!\bigl{[}\,Z_{k}^{(0)}=0\mid Z_{k}^{(1)}=1\,\bigr{]}.$
Feature Representation Confidence (FRC).
We combine the two causal probabilities with a harmonic mean to penalise vectors that are only sufficient or only necessary:
$\operatorname{FRC}_{k}=2\cdot\frac{\operatorname{PS}_{k}\,\operatorname{PN}_{k}}{\operatorname{PS}_{k}+\operatorname{PN}_{k}}.$
We first perform sensitivity pre-filtering by computing $\operatorname{EALE}_{k}$ for every base vector and retaining those whose absolute value exceeds the 75th percentile; on this reduced set we estimate $\operatorname{PS}_{k}$ and $\operatorname{PN}_{k}$ from every $\langle s^{+},s^{-}\rangle$ pair and rank the vectors by their $\operatorname{FRC}_{k}$; finally, the activation distributions of the top-10 ranked vectors are passed to an LLM agent, which verifies that each vector genuinely encodes the intended linguistic feature and flags any inconsistent or spurious patterns.
@@ 3.3 Feature Intervention
When we modify the values of SAE’s activation during forward propagation, we expect that such targeted interventions will influence the model’s behavior. However, our experiments show that altering only a small subset of features may not significantly impact the output—likely because linguistic phenomena are represented by multiple features across various layers. To assess the true impact of these interventions, we use a large language model as a judge. For each linguistic feature, we conduct both ablation and enhancement experiments. In the ablation experiment, we set the target feature’s activation to 0, and in the enhancement experiment, we set it to $10$. In both cases, we also perform baseline experiments by randomly selecting 25 base vectors from the same layer.
@@
Let $P_{\mathrm{abl}}^{T}$ and $P_{\mathrm{abl}}^{B}$ denote the success probabilities for the targeted and baseline ablation experiments, respectively. The normalized ablation effect is
$E_{\mathrm{abl}}=\frac{P\bigl{(}Y=0\mid I_{\mathrm{abl}}^{T}\bigr{)}-P\bigl{(}Y=0\mid I_{\mathrm{abl}}^{B}\bigr{)}}{P\bigl{(}Y=0\mid I_{\mathrm{abl}}^{T}\bigr{)}}.$
The normalized enhancement effect $E_{\mathrm{enh}}$ is defined analogously as the difference between targeted and baseline enhancement success probabilities, normalized by $1-P\bigl{(}Y=1\mid I_{\mathrm{enh}}^{B}\bigr{)}$.
@@
When calculating FIC, if one or both of the $E$ values are negative, we incorporate a penalty coefficient $w$ to reflect the weakened or lost causality in such cases. This FIC score provides a balanced measure of how effectively targeted interventions, as opposed to random ones, influence the model’s output with respect to specific linguistic features. The details for FIC are shown in Appendix E.2.